### v1

# MoCA: Multi-modal Cross-masked Autoencoder for Digital Health Measurements

Abstract
The growing prevalence of digital health technologies has led to the generation of complex multi-modal data, such as physical activity measurements simultaneously collected from various sensors of mobile and wearable devices. These data hold immense potential for advancing health studies, but current methods predominantly rely on supervised learning, requiring extensive labeled datasets that are often expensive or impractical to obtain, especially in clinical studies. To address this limitation, we propose a self-supervised learning framework called Multi-modal Cross-masked Autoencoder (MoCA) that leverages cross-modality masking and the Transformer autoencoder architecture to utilize both temporal correlations within modalities and cross-modal correlations between data streams. We also provide theoretical guarantees to support the effectiveness of the cross-modality masking scheme in MoCA. Comprehensive experiments and ablation studies demonstrate that our method outperforms existing approaches in both reconstruction and downstream tasks. We release open-source code for data processing, pre-training, and downstream tasks in the supplementary materials. This work highlights the transformative potential of self-supervised learning in digital health and multi-modal data.
Machine Learning, ICML
1 Introduction
Advancements in mobile and wearable device technology in recent years have enabled the measurement and tracking of complex multi-modal health data. For example, physical activity can be continuously monitored using multiple sensors, such as accelerometers and gyroscopes, embedded in wearable trackers. The scope of digital health data extends beyond activity tracking to modalities such as Electroencephalography (EEG) (Xie et al., 2022a; Yuan et al., 2024), sleep patterns (Fayyaz et al., 2023), glucose levels (Lee et al., 2023; Sergazinov et al., 2023), functional magnetic resonance imaging (fMRI) (Sarraf et al., 2023; Kwon et al., 2025), and many more. These rich data have become instrumental in disease prevention, early diagnosis, treatment monitoring, and intervention assessment, offering unprecedented opportunities to advance health research and improve patient outcomes.
Our work is motivated by the practical application of mobile and wearable trackers for monitoring physical activity through multiple sensors. This application has gained increasing popularity in clinical studies, enabling researchers to examine how physical activity influences various health outcomes such as cardiovascular disease, and to facilitate early interventions for conditions such as the Alzheimer’s disease. By understanding the role of physical activity in health, researchers and clinicians can also develop evidence-based guidelines for targeted interventions, ultimately promoting disease prevention and overall well-being.
In studies involving data from wearable and mobile devices, measurements are often recorded in “free-living” environments without laboratory calibration, resulting in a lack of gold-standard labels. Missing data from one or more modalities is also a common challenge. These issues are especially pronounced in clinical studies, where sample sizes tend to be small and collecting calibrated data is more difficult (Di et al., 2022). In such scenarios, self-supervised learning has emerged as an effective approach by leveraging unlabeled data in pre-training. A particularly promising approach within this paradigm is the masked autoencoder (MAE) (He et al., 2021), which was originally developed for computer vision but can be adapted for time series data (Li et al., 2023; Zhang et al., 2023; Tang & Zhang, 2022). In MAE, a portion of the input data is randomly masked, and the model is trained to reconstruct these masked elements using only the unmasked information to learn latent representations that capture essential patterns in the data.
Comparing to data from a single modality, data from multiple modalities have more complex correlation structures, including inter- and intra-modality correlations. In digital health applications, these correlations are often temporal in nature, with simultaneously measured signals across different modalities providing insights into each other. When developing multi-modal MAE approaches, it is crucial to take into account such correlation structures. While several MAE methods have been proposed for multi-modal data (Yang et al., 2024; Bachmann et al., 2022; Yan et al., 2022), existing approaches often lack theoretical justification for their masking strategies and instead rely on ad hoc solutions.
We propose the Multi-modal Cross-masked Autoencoder (MoCA), a novel approach that incorporates both intra- and inter-modality correlation structures in its masking strategy. Our theoretical analysis establishes guarantees for how different masking patterns influence the lower bound of the MAE loss function. Through a Reproducing Kernel Hilbert Space (RKHS) framework, we demonstrate a fundamental connection between multi-modality MAE loss and canonical correlation analysis (CCA). This connection not only explains MoCA’s superior performance compared to alternative approaches, but also suggests how correlation structures in multi-modal data can guide the development of optimal masking strategies.
We evaluate MoCA through extensive experiments and ablation studies using real-world physical activity data collected from mobile devices across multiple modalities. Our analysis examines both reconstruction accuracy and performance on downstream tasks. The results demonstrate that MoCA’s cross-modality masking scheme consistently outperforms existing baselines. Through detailed visualizations of attention maps in downstream tasks, we demonstrate how MoCA effectively generates informative latent representations to predict activity types.
2 Related Work
Self-supervised Learning (SSL) has emerged as a powerful framework to learn useful representations from unlabeled data in computer vision (Pathak et al., 2017; Gidaris et al., 2018; Wang & Gupta, 2015). SSL has been coupled with Transformer-based models (Xie et al., 2021; Li et al., 2021; Caron et al., 2021), and has been extended to time series data (Zerveas et al., 2020; Baevski et al., 2022; Luo et al., 2023; Nie et al., 2023; Xie et al., 2022b) and digital health (Krishnan et al., 2022), offering pathways to learning meaningful representations without extensive labeled datasets. SSL also forms the foundation of large language modeling (LLM) (Devlin et al., 2019)(Brown et al., 2020).
Masked Autoencoder (MAE) (He et al., 2021) is a type of Masked Image Modeling (MIM) self-supervised approach in vision, utilizing masked patches of images and autoencoder framework. Since its development, multiple works have provided theoretical understanding of MAE (Cao et al., 2022; Pan et al., 2022; Zhang et al., 2022b). As a vision model, MAE has been used in a wide variety of domains (Zhang et al., 2022a), such as with audio input (Gong et al., 2022; Huang et al., 2022), with video input (Feichtenhofer et al., 2022; Tong et al., 2022), and in time-series domain (Li et al., 2023; Zhang et al., 2023; Dong et al., 2023).
Multi-modality learning focuses on modeling information across multiple data modalities such as time series from multiple sensors, images, and text. The inherent challenge lies in capturing the complex inter-dependencies between modalities. Representative works in SSL with multi-modalities features include (Singh et al., 2022; Radford et al., 2021; Yu et al., 2021; Taleb et al., 2021), MAE with multi-modality spans vision and language (Geng et al., 2022), wearable device data (Liu & Liu, 2024), audio-visual (Gong et al., 2022), and with missing modalities (Liu et al., 2023b). Representative large vision language multi-modality model also include (Liu et al., 2023a) and (Alayrac et al., 2022). Existing approaches often rely on ad hoc masking strategies without theoretical justification. In this work, we develop a multi-modal MAE approach specifically designed for digital health data and, to our knowledge, provide the first theoretical guarantee for the reconstruction bound of multi-modal MAE under different masking schemes.
3 Proposed Method
We propose the Multi-modal Cross-masked Autoencoder (MoCA), a masked autoencoding approach designed specifically for multi-modal data, with a focus on time series measurements (for detailed model architecture see Appendix B). MoCA accounts for intra- and inter-modality correlation structures through specialized tokenization and masking procedures as the following.
3.1 Tokenization
[…]
Channel-independent tokenization independently projects time windows of different modalities into different tokens (patches). Input $x$ is first pachified into $\mathbb{R}^{C\times P\times L_{p}}$. This is then flattend to 1D sequence of time windows, resulting in $x\in\mathbb{R}^{(C\cdot P)\times L_{p}}$. A linear transformation $\mathbb{R}^{(C\cdot P)\times L_{p}}\mapsto\mathbb{R}^{(C\cdot P)\times D}$ follows to project $CP$ tokens to dimension $D$, where $D$ is the embedding dimension for the encoder. A fixed 2D sinusoidal positional embedding is added to retain both temporal and modality information.
3.2 Masking Schemes
With a specified mask ratio $\rho$, we randomly mask a proportion $\rho$ of the tokenized time series patches. The model then uses a Vision Transformer (ViT) (Dosovitskiy et al., 2021) architecture, consisting of an encoder and decoder (detailed in Appendix B), where the encoder processes the unmasked patches to generate latent representations, which the decoder uses to reconstruct the masked patches. The choice of which patches to mask is crucial. In MoCA, we propose the following masking schemes:
Cross-modality masking. This approach masks patches in different modalities independently across time windows. Intuitively, cross-modality masking introduces more flexibility, leading to a more diverse distribution of modalities at each time window in both unmasked and masked patches. In multi-modal health data, patches from different modalities often exhibit strong cross-modality correlations in the same and nearby time windows. The proposed cross-modality masking scheme leverages cross-modal correlations effectively to enhance representation learning between unmasked and masked patches.
Synchronized masking. As a comparison, an alternative approach is to simultaneously mask patches from all modalities at the same time windows. Several MAE approaches for multivariate time series adopt the synchronized masking (Li et al., 2023; Zhang et al., 2023) by treating multivariate time series from different modalities as channels when applying the vision MAE. Essentially, synchronized masking is a constrained masking strategy in which patches from different modalities must be selected as masked or unmasked simultaneously within the same time window.
3.3 Data Augmentation
To further enhance MoCA’s cross-modality learning capabilities, we augment our dataset by incorporating more time series samples containing state transitions. Specifically, we use a straightforward augmentation strategy that can be applied both offline and on-the-fly. In this approach, two time series samples are randomly selected, and a mixup operation (Zhang et al., 2018) is performed. A segment of variable length (between 20% and 50% of the original time series length) is extracted from one sample and inserted into the corresponding position of the other. This process is repeated to generate the desired number of augmented samples and saved to disk for reproducibility. Further details about augmentation are provided in Appendix D Algorithm 1.
4 Theoretical Results
We present here theoretical results on the bounds of the MAE loss function under different random masking schemes. Additionally, we establish a connection between the MAE loss function bounds and canonical correlation analysis (CCA) using a Reproducing Kernel Hilbert Space (RKHS) approach.
4.1 Notation
Let $\bar{x}$ denote a generalized “image” from a dataset that consists of a matrix of data patches of multiple modalities. Each row of $\bar{x}$ corresponds to one modality. Given a total of $C$ modalities, suppose each modality contains $P$ patches, where each patch is a time series of a fixed length. Then, each image consists of a total of $CP$ patches.
[…]
Given a mask ratio $\rho$, a realization of the random masking matrix $\delta$ is a matrix taking values of 0 or 1, where the total number of ones is $\lfloor\rho CP\rfloor$. Conditional on $\bar{x}$, the probability distribution of $\delta$ over all admissible masking matrices, subject to the constraints of the sampling scheme, determines the distribution of masked and unmasked views.
4.2 Bounds for MAE Reconstruction Losses
The Masked Autoencoder (MAE) model consists of an encoder $f$, which maps the unmasked view $x^{U}$ to a latent feature $z^{U}=f(x^{U})$, and a decoder $g$, which maps $z^{U}$ back to the space of the masked view $x^{M}$. The objective of the MAE is to minimize the loss function
$L_{\text{MAE}}=\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\|g\circ f(x^{U})-x^{M}\|^{2}.$
We assume that $x^{M}$, $x^{U}$, and the reconstructed masked view $\hat{x}^{M}:=g\circ f(x^{U})$ are all $l_{2}$-normalized to the same scale.
The following assumption is from (Zhang et al., 2022b):
Assumption 4.1.
For any non-degenerate decoder $g$, assume there exists a pseudo-inverse encoder $f_{g}$ such that the pseudo autoencoder $h_{g}:=g\circ f_{g}$ satisfies $\mathbb{E}_{x}\|h_{g}(x)-x\|^{2}\leq\epsilon$, where $x$ represents either unmasked view $x^{U}$ or masked view $x^{M}$ and $\epsilon$ is a small positive constant.
It has been shown in (Zhang et al., 2022b) that given Assumption 4.1, the MAE loss is bounded from below:
$L_{\text{MAE}}(h)\geq-\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle h(x^{U}),h_{g}(x^{M})\rangle-\epsilon+\text{const},$ (1)
[…]
4.3 Kernel Representations
We represent the encoder and decoder mappings using Reproducing Kernel Hilbert Space (RKHS) representations. Assume there exist mappings $\Phi$ and $\tilde{\Phi}$ such that
$h(x^{U})=g\circ f(x^{U})=[\Phi(x_{c_{i_{1}},p_{i_{1}}}),\dots,\Phi(x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})],$ (2)
where $\{i_{1},\dots,i_{n_{u}}\}$ are indices of the unmasked patches, and $n_{u}=CP-\lfloor\rho CP\rfloor$ is the total number of patches in the unmasked view $x^{U}$, and similarly,
$h_{g}(x^{M})=g\circ f_{g}(x^{M})=[\tilde{\Phi}(x_{c_{j_{1}},p_{j_{1}}}),\dots,\tilde{\Phi}(x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})],$ (3)
where $\{j_{1},\dots,j_{n_{m}}\}$ are indices of the masked patches, and $n_{m}=\lfloor\rho CP\rfloor$ is the total number of patches in the masked view $x^{M}$.
Using the kernel trick, we can further write (2) using the representation
$\Phi(x)=\sum_{k}\alpha_{u}^{T}(k)k_{u}(x,x_{c_{i_{k}},p_{i_{k}}}),$
where $k_{u}(\cdot,\cdot)$ is the kernel associated with the RKHS in which $\Phi$ resides in, so that for all pairs of unmasked patches $(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})$,
$k_{u}(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})=\langle\Phi(x_{c_{i_{k}},p_{i_{k}}}),\Phi(x_{c_{i_{l}},p_{i_{l}}})\rangle,$
and $\alpha_{u}$ is a vector ($\alpha^{T}_{u}$ is its transpose) in which the $k$th element $\alpha_{u}(k)$ is the weight of the basis spanned by the reproducing kernel at unmasked patch $x_{c_{i_{k}},p_{i_{k}}}$.
Similarly,
$\tilde{\Phi}(x)=\sum_{l}\alpha_{m}^{T}(l)k_{m}(x,x_{c_{j_{l}},p_{j_{l}}}),$ (4)
where $k_{m}(\cdot,\cdot)$ is the kernel associated with the RKHS in which $\tilde{\Phi}$ resides in, so that for all pairs of masked patches, $(x_{c_{j_{k}},p_{j_{k}}},x_{c_{j_{l}},p_{j_{l}}})$,
$k_{m}(x_{c_{j_{k}},p_{j_{k}}},x_{c_{j_{l}},p_{j_{l}}})=\langle\tilde{\Phi}(x_{c_{j_{k}},p_{j_{k}}}),\tilde{\Phi}(x_{c_{j_{l}},p_{j_{l}}})\rangle,$
and $\alpha_{m}$ is a vector, in which the $l$th element $\alpha_{m}(l)$ is the weights of the basis spanned by the reproducing kernel at masked patch $x_{c_{j_{l}},p_{j_{l}}}$.
4.4 Connection to Canonical Correlation Analysis (CCA)
Minimizing the lower bound in (1) is equivalent to maximizing the quantity $\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle h(x^{U}),h_{g}(x^{M})\rangle$, which, using the RKHS representations in the previous section, can be written as
$\begin{split}\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle[\Phi(x_{c_{i_{1}},p_{i_{1}}}),\dots,\Phi(x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})],\\ [\tilde{\Phi}(x_{c_{j_{1}},p_{j_{1}}}),\dots,\tilde{\Phi}(x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})]\rangle,\end{split}$
which equals to
$\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}[\alpha_{u}^{T}K_{u}K_{m}\alpha_{m}],$ (5)
where $K_{u}=k_{u}-k_{u}\mathbf{1}-\mathbf{1}k_{u}+\mathbf{1}k_{u}\mathbf{1}$ with $k_{u}$ denoting the matrix whose $(k,l)$th element is $k_{u}(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})$ and $\mathbf{1}$ denotes the matrix with all elements equal to one. The definition of $K_{m}$ follows the same pattern.
The representation of (5) has a close connection to the Canonical Correlation Analysis (CCA). Given a distribution of random images ${\bar{x}}$, the random masking scheme, characterized by the distribution of $\delta$, determines the distribution of unmasked and masked view assignments. For each pair $(x^{U},x^{M})$, the matrices $k_{u}$ and $k_{m}$ are determined by their respective compositions $x^{U}=(x_{c_{i_{1}},p_{i_{1}}},\dots,x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})$ and $x^{M}=(x_{c_{j_{1}},p_{j_{1}}},\dots,x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})$.
Given a realization of $\delta$, the resulting $(x^{U},x^{M})$, and under the assumption of $l_{2}$ regularization of both unmasked and masked views, the optimal solution to (5) is obtained by finding vectors $\alpha^{*}_{u}$ and $\alpha^{*}_{m}$ that satisfy
$(\alpha^{*}_{u},\alpha^{*}_{m})=\operatorname*{arg\,max}_{\alpha_{u},\alpha_{m}}\frac{\alpha^{T}_{u}K_{u}K_{m}\alpha_{m}}{\sqrt{(\alpha^{T}_{u}K_{u}^{2}\alpha_{u})(\alpha^{T}_{m}K_{m}^{2}\alpha_{m})}},$ (6)
which can be solved using the Lagrangian multiplier (for details, see, for example, (Yang et al., 2019; Andrew et al., 2013)).
A special case is when the mappings $\Phi$ and $\hat{\Phi}$ are linear, then the objective in (6) becomes the classical CCA objective
$\max_{\alpha_{u},\alpha_{m}}\frac{\alpha^{T}_{u}\Sigma_{UM}\alpha_{m}}{\sqrt{(\alpha^{T}_{u}\Sigma_{UU}\alpha_{u})(\alpha^{T}_{m}\Sigma_{MM}\alpha_{m})}},$ (7)
where $\Sigma_{UM}$ denotes the cross covariance matrix between all pairs of unmasked and masked patches, $\Sigma_{UU}$ denotes the cross covariance matrix of all pairs of unmasked patches, and $\Sigma_{MM}$ denotes the cross covariance matrix of all pairs of masked patches. The solution to this problem will be discussed in the next section.
Figure 1: Example image in the simulation study. $x_{c,p}$ denotes $p$th patch in $c$th modality. Note how patches from different modalities within the same time window exhibit strong cross-modality correlations.
4.5 Influence of Masking Schemes
Since any random masking scheme defines a probability measure over the space of admissible masking realizations of $\delta$, the objective in (5) can be interpreted as a weighted average of the optimal values of the objective in (6). If the matrices $K_{u}$ and $K_{m}$ are known or can be reasonably approximated, then the optimal random masking scheme is one that assigns higher probabilities to masking allocations that yield larger objective values.
In the special case of linear kernels, the optimal objective is the top singular value of the matrix $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ (Sánchez et al., 1979; Andrew et al., 2013). Therefore, to optimize (5), higher probabilities should be assigned, subject to the mask ratio constraint, to masking allocations that result in matrices $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ with larger top singular values.
The matrix $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ can be interpreted as a normalized adjacency matrix of a bipartite graph, where nodes are masked and unmasked patches, and edge weights are values in the corresponding elements of the matrix. A large top singular value indicates a potentially strong community/cluster of nodes that are highly correlated with each other (Sarkar & Dong, 2011), that is, pairs of masked and unmasked patches with large absolute values of the correlation. Intuitively, to optimize (5), the random masking scheme should generate such pairs with larger probabilities.
4.6 Demonstration with Synthetic Data
To validate the above heuristic, we compare MoCA against synchronized masking through simulations with synthetic data. Following Algorithm 1, we generate synthetic data by sampling two samples $\bar{x}_{1},\bar{x}_{2}$ from the training set, and then randomly extract a consecutive chunk of time series from $\bar{x}_{2}$ to replace the corresponding time series in $\bar{x}_{1}$. We repeated this process 1000 times to obtain our simulation data. An example image is shown in Figure 1. With both MoCA and synchronized masking, we calculate the averaged optimal objective in (6). MoCA yields a value of 0.1989, comparing to 0.1812 with synchronized masking. This result demonstrates that MoCA generates masked and unmasked patches that better capture cross-modal correlations, leading to improved reconstruction, whereas synchronized masking fails to leverage correlations between patches from different modalities within the same time window.
5 Experiments
In this section, we present experimental justifications to our hypothesis, and compare MoCA with the state-of-the-art self-supervised learning approaches. We follow the evaluation protocol in (He et al., 2021), by pre-training the network using MoCA, and compare the generalization ability of the obtained features out of the network. We pre-train the ViT-Base (Dosovitskiy et al., 2021) model on the UCI-HAR training set (Reyes-Ortiz et al., 2013). Data augmentation was applied to MoCA, which significantly scales the dataset from 1,847 to 369,400 samples.
We consider the following the tasks for our model evaluation: 1) the linear probing (Chen et al., 2021) and finetuning accuracy on the classification task using the frozen pre-trained network parameters, 2) the transfer learning task by initialization from the pre-trained network parameters, and 3) the prediction accuracy on generative tasks. For more details on the UCI-HAR dataset, see Appendix A.
5.1 Pre-training Set-ups and Optimization
We pre-train MoCA on ViT-Base (architecture details in Appendix B) for 4,000 epochs using the AdamW optimizer with an initial learning rate of 0.0005 and a batch size of 50 (See Appendix D for implementation details). To ensure a fair comparison, we limited MoCA’s training on the augmented datasets (described in Section 3.3) to 20 epochs, aligning the total number of training iterations. We train the methods either with (w/) or without (w/o) augmentation (aug.) during training. See Appendix D, Algorithm 1 for more details on how augmentations are generated.
One of our most critical baseline is MAE with ViT-Base (He et al., 2021) with synchronized masking scheme across modalities (abbreviated as “MAE Synchronized Masking”) pre-trained on UCI-HAR, where all modalities are masked or unmasked altogether at given time windows (i.e., synchronized masking). The MAE Synchronized Masking has an input size of [6, 200 1], corresponding to [row, column, channel], since we concatenate the 6 channels into 1 image as rows, and randomly sample on each column at the same time according to mask ratio.
We also compare MoCA with MAE with ViT-Base pre-trained on ImageNet1k (Russakovsky et al., 2015) (“MAE-Base w/ IN1k”) and BEiT v2 (Peng et al., 2022) pre-trained on UCI-HAR with code-book obtained from ImageNet-1k.
Table 1: Classification performance of self-supervised models in finetuning (FT) and linear probing (LP) tasks on the UCI-HAR dataset. Top 1 accuracies (%) for FT and LP are reported.
Method FT Top-1 (%) LP Top-1 (%)
ViT-Base (Dosovitskiy et al., 2021) 91.4 /
MAE-Base w/ IN1k (He et al., 2021) 74.0 86.5
BEiT v2 (Peng et al., 2022) 70.4 86.1
MAE Synchronized Masking (w/o aug.) 91.7 86.3
MAE Synchronized Masking (w/ aug.) 91.3 88.4
MoCA (w/o aug., Ours) 94.8 92.4
MoCA (w/ aug., Ours) 96.6 93.1
5.2 Classification Task
Table 2: Limited labels top-1 accuracy for finetuning (FT) and linear probing (LP) on UCI-HAR. MAE Synchronized Masking and MoCA without data augmentation are compared in terms of $1\%$ and $10\%$ limited labels.
Model Labels FT Top-1 (%) LP Top-1 (%)
MAE Synchronized Masking 1% 57.8 52.5
10% 84.2 77.7
100% 91.7 86.3
MoCA (Ours) 1% 76.0 67.3
10% 87.1 86.4
100% 94.8 92.4
Table 3: Transfer learning performance of MoCA on unseen activity recognition datasets. Top 1 accuracies (%) for finetuning (FT) and linear probing (LP) are reported.
Model WISDM IM-WSHA
FT Top-1 (%) LP Top-1 (%) FT Top-1 (%) LP Top-1 (%)
MAE-Base w/ IN1k (He et al., 2021) 77.91 86.92 36.51 60.32
MAE Synchronized Masking (w/o aug.) 92.73 87.79 66.67 58.73
MAE Synchronized Masking (w/ aug.) 92.73 88.08 60.32 60.32
MoCA (w/o aug., Ours) 93.02 90.70 57.14 52.38
MoCA (w/ aug., Ours) 93.90 93.60 74.60 71.43
Linear Probing. We evaluate methods on the UCI-HAR test set by following the linear probing protocol (He et al., 2021), where the pre-trained parameters are frozen, and only the attached classifier is trained during downstream task training (see Appendix D for implementation details). We report top-1 accuracy on the UCI-HAR test set. We compare MoCA with 1) MAE-Base w/ IN1k, 2) MAE Synchronized Masking, and 3) BEiT v2 (Peng et al., 2022). As shown in Table 1, the distribution gap between time series and images results in poor performance for MAE-Base w/ IN1k and BEiT v2. Both MoCA w/ aug. and MoCA w/o aug. have surpassed the baselines by significant margins, as cross-modality time-series patches embedded important sequential dependencies among patches. The spatiotemporal cross-modality synchronization enables the network to learn more semantic correlation across modalities.
Finetuning. We finetune the pre-trained models using the labels provided in the UCI-HAR training set, and compare the classification accuracy on the UCI-HAR test set. MoCA surpasses the MAE Synchronized Masking by a margin over $3\%$ point, thanks to its exclusive spatiotemporal mask encoding scheme. The augmentation has also effectively improved the generalization of MoCA, when different modalities are jointly learned in an non-synchronized way.
Semi-supervised learning. In this section, we test the model performances when using limited labels for downstream classification, following the training protocol in (Caron et al., 2020). Finetuning and linear probing were conducted with $1\%$ and $10\%$ labels of UCI-HAR based on pre-trained MoCA. The $1\%$ and $10\%$ limited samples are representative (keeps the same class ratios) of the overall activity label distribution from the entire training set. For simplicity, we only compare MoCA and MAE Synchronized Masking w/o data augmentation. Table 2 shows the top-1 finetuning and linear probing for 7-class classification accuracy. As can be seen, MoCA demonstrates superior performance against MAE Synchronized Masking baseline. The performance gap between MoCA and the MAE Synchronized Masking baseline is emphasized when the model is finetuned with only $1\%$ labels. This supports our hypothesis that when very limited number of labels are available during downstream tasks, the proposed cross-modality MoCA sampling scheme provides better initialization and generalization.
Supervised learning. To evaluate the effectiveness of MoCA sampling scheme, we also trained a ViT-Base from scratch on the UCI-HAR dataset for 50 epochs, which has the same training iteration of finetuning MoCA. The linear probing accuracy of MoCA surpassed that of supervised learning methods, demonstrating its capacity to generalize by leveraging on effective pre-training. Notably, finetuning ViT with MoCA’s pre-trained weights yielded a significant accuracy increase from 91.4% to 96.6% (Table 1).
5.3 Transfer Learning
One solution to address time series classification task is to view multivariate time series as 2D images, and to use vision encoders with pre-trained weights from ImageNet to yield prediction (Wimmer & Rekabsaz, 2023; Vishnupriya & Meenakshi, 2018; Chun et al., 2016; Krishnan et al., 2020). However, due to the substantial differences between image data and activity signals, it remains doubtful whether vanilla out-of-domain image pre-training on ImageNet is beneficial for learning useful time series representations. To investigate this, we compare the transferability of MoCA against the ViT-Base model pre-trained on ImageNet weights via vanilla masked autoencoders (MAE) training, without any additional finetuning. We evaluate MoCA on two unseen datasets: WISDM (Kwapisz et al., 2011) and IM-WSHA (Tahir et al., 2020; Jalal et al., 2020). WISDM, which contains only accelerometer data and six activities, tests MoCA’s ability to handle incomplete data. IM-WSHA includes both accelerometer and gyroscope data with 11 activities, allowing for a more comprehensive assessment. Downstream training details are in Appendix D, and datasets details are in Appendix A.
Figure 2: Evaluation of imputation tasks on UCI-HAR test set. The figure shows the results of three imputation tasks: random imputation (top), temporal imputation (middle), and sensor imputation (bottom). For both random and temporal imputation, 70% of the data is masked, while for sensor imputation, only a single channel/sensor remains visible. A single representative sample is randomly selected for each task for visualization.
Our empirical evidence, as presented in Table 3, indicates that pre-training on the conventional ImageNet1K alone is insufficient for effective time series data downstream tasks on different dataset. Specifically, it is notable that the performance of the finetuned out-of-distribution (OOD) pre-trained model (MAE-Base IN1K model) drops drastically even compared to its linear probing from the same pre-trained model. One explanation for this would be that finetuning in OOD settings distorts the pre-trained features, imposing challenges on improving downstream performance (Kumar et al., 2022). This supports the motivation that on-the-shelf ImageNet pre-trained model for time-series activity data is not effective.
Our proposed MoCA w/ aug., in contrast, surpasses the baselines in most linear probing and finetuning evaluations. Notably, it demonstrates robust adaptability to incomplete modality settings. We observed that on the WISDM dataset, MoCA requires only a few epochs of finetuning to transition from six-channel pre-training to a three-channel configuration, achieving competitive results using solely accelerometer data while excluding gyroscope inputs. This adaptability addresses a critical challenge in real-world scenarios, such as clinical practice, where incomplete sensor data is a common occurrence. Table 3 highlights that MoCA w/ aug. demonstrates better generalization on IM-WSHA compared to synchronized masking MAE, especially in scenarios where subject heterogeneity makes it challenging for the model to rely directly on temporal information. However, we also noticed that when there are limited pre-training data without augmentations, both the accuracies of MAE Synchronized Masking (w/o aug.) and MoCA (w/o aug.) drop drastically. As recognized by the community (He et al., 2021), this verifies the critical role of the large dataset in self-supervised learning, and justifies the effectiveness of our proposed augmentation generation tailored for digital health data.
5.4 Generative Task
Table 4: Performance comparison of imputation methods on the UCI-HAR test set. We measure the reconstruction MAE (Mean Absolute Error) and MSE (Mean Squared Error), where lower values indicate better performance.
Model / Error (MAE / MSE) $\downarrow$ Random Imputation Temporal Imputation Sensor Imputation Temporal Extrapolation
Linear Interp. 0.587 0.689 0.546 0.634 0.252 0.189 0.415 0.474
Nearest Neighbor 0.341 0.284 0.240 0.180 0.412 0.377 0.210 0.158
MICE 0.289 0.251 0.410 0.375 0.210 0.158 0.210 0.158
MAE Synchronized Masking (w/o aug.) 0.275 0.162 0.262 0.191 0.288 0.170 0.262 0.191
MAE Synchronized Masking (w/ aug.) 0.380 0.281 0.306 0.231 0.494 0.459 0.175 0.118
MoCA (w/o aug., Ours) 0.086 0.043 0.112 0.059 0.160 0.089 0.112 0.060
MoCA (w/ aug., Ours) 0.076 0.027 0.153 0.102 0.140 0.063 0.157 0.109
Table 5: Synchronized masking input size ablation results. Finetuning (FT) and linear probing (LP) top-1 accuracies on UCI-HAR are reported.
Model (input size) FT Top-1 (%) LP Top-1 (%)
MAE ([1, 200, 6]) 92.7 90.4
MAE Stacked inputs ([6, 200, 6]) 85.6 50.8
MAE Synchronized Masking ([6, 200, 1]) 91.7 86.3
MoCA (Ours) ([6, 200, 1]) 94.8 92.4
Wearable devices have transformed health monitoring and other domains, but missing data remains a persistent challenge that limits their reliability. Data gaps can arise from signal glitches, battery depletion, or device malfunction. These challenges are magnified when multiple sensors are used. In this section, we evaluate each methods with respect to real world challenges in health monitoring and assess each model’s ability to reconstruct missing data effectively. The tasks include the following: (i) Random Imputation tests the model’s ability to reconstruct localized missing values by randomly masking small data patches and predicting their original values, (ii) Temporal Imputation evaluates long-term dependency modeling by masking all sensor data over consecutive time periods and requiring the model to predict the missing segments, and (iii) Sensor Imputation measures cross-modality learning by masking an entire sensor channel and having the model reconstruct it using information from correlated signals.
Similar to (Narayanswamy et al., 2024), we will evaluate the reconstruction results on the UCI-HAR testing set against two widely used statistical imputation methods: (1) Linear Interpolation, which fills missing values by linearly interpolating between known values along the temporal dimension, and (2) Nearest Neighbor Imputation, which imputes missing values by averaging the values of the k-nearest neighbors to handle missing data. To evaluate the imputation methods, we use Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better performance.
Table 5 shows that MoCA (w/ aug.) significantly outperforms the traditional statistical imputation method and MAE Synchronized Masking baseline. However, in the temporal imputation task, MoCA (w/ aug.) performs worse than MoCA (w/o aug.), likely due to a domain shift from training on large-scale generated synthetic data. While augmentation can likely improve generalization, the feature might shift from reflecting authentic temporal dependencies in the downstream tasks.
Figure 2 indicates that cross-modality MAE is especially effective for improving imputation accuracy in multi-sensor data. In the random imputation task, MoCA uses only a few visible patches scattered across different channels to faithfully reconstruct the entire series. By drawing on cross-modality relationships, it captures shared information and reconstructs missing signals more accurately than other approaches. Moreover, during sensor imputation when all sensors except gyroscope-y (Gyr_Y) are masked, other methods collapse to predicting zero for all time steps. MoCA leverages the remaining signals to reconstruct realistic patterns across all sensors. This shows its ability to understand complex inter-sensor relationships, which is valuable for robust multi-sensor data analysis.
Figure 3: Attention maps from MoCA for predicting transition (top) and walking up (bottom). The patch colors reflect the attention scores, with yellow indicating low attention and dark red indicating high attention, where highly attended patches contribute more to the final prediction. The black lines illustrate the raw signal series. Patch size is 20.
5.5 Ablation Studies and Analysis
We examine the impact of input size changes upon MoCA’s concatenation of different channels. We illustrate the dimensions as [row, column, channel] as discussed in Section 3. We compare MoCA pre-trained on UCI-HAR with: 1) vanilla MAE with original input size [1, 200, 6], 2) vanilla MAE with stacked input where 6 sequence are concatenated into 6 rows of the image with input size [6, 200, 6], and 3) MAE Synchronized Masking where we concatenate the 6 channels into 1 image as rows, and randomly sample on each column at the same time according to mask ratio, input size [6, 200 1]. Training epoch is adjusted so that the total training iteration remains the same after adjusting for decrease in sample size for the stacked input. Regardless of how we adjust the size for vanilla MAE, the MoCA cross-modality sampling scheme always performs the best, showing that MoCA does not simply benefit from the change of input size, but instead essentially learned advantageous cross-modality information from the pre-training and the masking scheme. For ablations on other hyperparamters, see Appendix C.
We extract the attention map from the final layer of our model’s encoder, where an attention map represents the distribution of attention weights that indicate how much focus the model assigns to different input regions. Specifically, we aggregate the attention weights from the CLS token for each modality to its corresponding patches and average them across all heads. This process allows us to visualize the attention scores in Figure 3, for all patch tokens in each modality that attends to the CLS token that ultimately contributes to the final prediction after applying a linear classifier. Figure 3 has shed light on how our model actually performs classifications. We observe that MoCA model tends to focus on critical transition points when the object is changing states, while also attending to important periodic patterns for activities.
6 Conclusion
In this paper, we propose a self-supervised learning framework called Multi-modal Cross-masked Autoencoder (MoCA), specifically designed to facilitate self-supervised learning for digital health data and to address the challenge of limited annotations in healthcare AI applications. MoCA explicitly incorporates cross-modality masking within a Transformer-based autoencoder architecture to effectively capture temporal dependencies within and across modalities. Theoretical analysis supports our hypothesis that this cross-modality masking scheme enhances the model’s ability to generalize features, leading to improved representation learning. Empirical results further validate the superior performance of MoCA in various real-world digital health prediction tasks.
Impact Statement
This work presents novel application of self-supervised learning in digital health which helps with data scarcity. Our model has wide implications on the digital health sphere, and helps to overcome the limited label issue and privacy issue in the health data AI community.
However, self-supervised learning can be implemented by anyone who has no access to the true label of the data, and this might pose threats to the safety of the deployed model. Once our model is published and made public, such self-supervised model without proper fintuning or professional interpretations may possibly be misleading in its prediction and implications. We recommend careful and thoughtful application and reproduction of our methods.
Appendix A Dataset
UCI-HAR(Reyes-Ortiz et al., 2013) contains activity information of 30 subjects who performed a series of 6 activities (walking, walking upstairs, walking downstairs, sitting, standing, laying) in the same order. UCI-HAR includes inertial measurement unit (IMU) signals in 6 modalities in total (accelerometer and gyroscope readings in x, y, and z axes) collected by a waist-worn smartphone at 50Hz. We segment each time series using a 4-second sliding window. If a window contains two activity labels and each activity lasts more than 1.2 seconds, we assign the segment the label “transition.” This approach creates a more challenging classification task with 7 classes.
In the pre-training stage, each sub-sequence has corresponding activity as the target for the downstream tasks, with the transition label indicating that more than two activities were captured in a sub-sequence. No further preprocessing including normalization is added. Training and test data split was done with 7:3 ratio.
WISDM(Kwapisz et al., 2011) consists of data sampled at 20Hz from 51 subjects, who performed 6 activities. Examples of these activities include walking, jogging, walking upstairs, walking downstairs, sitting and standing. The dataset only includes measurements from the accelerometer embedded in a smartphone, thus our analysis will primarily focus on the robustness of our pre-trained model against partially missing modalities.
IM-WSHA(Tahir et al., 2020; Jalal et al., 2020) includes motion data sampled at 100Hz from three triaxial inertial measurement unit (IMU) sensors placed on the wrist, chest, and thigh of subjects. Participants performed 11 distinct activities, including computer use, phone conversation, vacuuming, reading, watching TV, ironing, walking, exercising, cooking, drinking, and brushing hair. For our analysis, we focus on the wrist IMU, which includes 220 sequences of inertial data, each lasting between 45 and 60 seconds.
Appendix B Model Architecture
Figure 4: Our MoCA architecture. During training, each time series is tokenized using a non-overlapping sliding window. A large portion of the tokens is then masked ($75\%$). The encoder processes only the observed tokens, while a small decoder operates on both the processed and masked tokens to reconstruct the input series.
The proposed model is based on the MAE proposed by (He et al., 2021) with some adaptations on the patch embedding adjusting for the multivariate time series input. MAE, as an autoencoder architecture, has an encoder and decoder sequence through which the input is projected to the learned representations (encoder), and reconstructed (decoder). MAE has Vision Transformer (ViT) (Dosovitskiy et al., 2021) encoder and decoder with masked patches. Downstream classification task involves finetuning and linear probing based on the pre-trained model.
Time series input. We have as our input a multivariate time series which warrants different pre-processing and model specification from the original MAE model. The MAE model supposes a collection of samples $x\in\mathbb{R}^{H\times W\times C}$ where $H$ and $W$ refer to the height and width of the image respectively with $C$ denoting number of channels. In the case of time series application, it is usually the case that $x\in\mathbb{R}^{1\times L\times C}$, where $L$ is the length of the time series with $C$ representing different modalities.
Encoder. MoCA uses ViT-Base (ViT-B), a stack of 12-layer standard Transformer blocks (Vaswani et al., 2023), as its encoder. Only the $\left\lfloor 0.25P\right\rfloor$ non-masked patches are fed into the Transformer blocks.
Decoder. The decoder is also composed of standard Transformer blocks of fewer layers, following the asymmetric architecture of MAE (He et al., 2021) of depth 8. The encoded patches are first padded with trainable masked tokens in its original position. Then, positional embeddings are added to this set of tokens, which give them information of their respective location. At the top of the decoder stack, we add a linear head to predict and reconstruct the input time series.
Objective. The MAE decoder learns to reconstruct the input multivariate series by predicting the values in each data point. The objective is to reduce the loss in the form of mean squared error (MSE) between the prediction and the input multivariate time series, averaged over all patches. This differs from the original implementation of the MAE model in that in (He et al., 2021), the loss is with respect to only the unmasked patches. This is to have a pre-trained model that has more comprehensive reconstruction performance across all the patches.
Appendix C Ablation Study
Additional ablation studies are conducted with regards to different mask ratios, patch lengths, and input lengths to investigate their effects on downstream classification top-1 accuracy. All ablation studies are conducted on MoCA without augmentation to ensure a fair comparison. Figure 5 summarizes the ablation experiments result.
Figure 5: MoCA ablation study results. Finetuning and linear probing top-1 accuracy on UCI-HAR is presented with different levels of mask ratio, patch size, and input length. The ablation study is performed with MoCA with ViT-Base and 4,000 training epochs.
Mask ratio. Using masking ratio from 5% to 95% with 5% increments, MoCA shows the best top-1 accuracy performance for mask ratio 75% or 55% for finetuning at 94.8%, and 55% for linear probing at 93.9%. This is mostly consistent with the MAE model originally suggested by (He et al., 2021), where mask ratio 75% was suggested for the model with image as inputs. Top-1 accuracy drops down quickly after 75% for both finetuning and linear probing.
Patch Length. The experiment result shows the effect of different patch length ($L_{p}$) when $L_{p}=\{5,10,20,25,40,50\}$. With the input length 200, this is equivalent to having 240, 120, 60, 48, 30, and 24 patches per input respectively. Patch length 5, 10 and 20 shows higher top-1 accuracies at 94.80%, 94.50%, and 94.80% respectively for finetuning and 93.30%, 92.80%, and 92.40% respectively for linear probing. Intuitively, having larger patch size means a larger chunk of input data at a time given a fixed mark ratio, which equates less overall exposure to the input data.
Input Length. We differed the input length ($L$) where $L=\{100,200,400,800,1600\}$. The number of patches per one input series stayed fixed at 10 across different input length, making the corresponding patch length $L_{p}=\{10,20,40,80,160\}$ respectively. Input length $L=200$ shows the best classification performance. Larger input lengths produced lower top-1 accuracies. Fixing the number of patches leads to inputs with larger lengths having longer patch lengths which results in lower accuracies. This coincides with what is generally suggested by the patch length ablation studies.
Appendix D Implementation Details
The hyperparameter configurations for pre-training, finetuning, and linear probing are detailed in Tables 8, 8, and 8, respectively, following established practices and adapted from Meta’s codebase (Touvron et al., 2021).
Table 8 summarizes the hyperparameters used for pre-training. It specifies that AdamW serves as the optimizer, with an absolute learning rate of 5e-4 and a weight decay of 0.05. The momentum parameters are $\beta_{1}$=0.9 and $\beta_{2}$=0.95, and the batch size is 50. A cosine decay learning rate schedule is employed, coupled with 50 warmup epochs. The model is trained for 20 epochs if data augmentation is applied, or 4000 epochs if no augmentation is applied.
Table 8 outlines the end-to-end finetuning settings for three datasets: UCI-HAR, WISDM, and IM-WSHA. All configurations use AdamW with absolute learning rate of 1e-3 for UCI-HAR, base learning rate of 1e-2 for WISDM, and base learning rate of 1e-3 for IM-WSHA. The weight decay is set to 0.05, and the momentum parameters are $\beta_{1}$=0.9, $\beta_{2}$=0.999. The batch size is 50 for UCI-HAR and WISDM, and 64 for IM-WSHA. A cosine decay schedule is used, with a warmup period of 5 epochs and a total of 50 training epochs.
Table 8 shows the hyperparameters for linear probing on these datasets. The AdamW optimizer is again adopted, but the weight decay is reduced to 0. The learning rate is absolute learning rate of 1e-3 for UCI-HAR, base learning rate of 1e-1 for WISDM, and base learning rate of 1e-2 for IM-WSHA, with the same momentum settings ($\beta_{1}$=0.9, $\beta_{2}$=0.999). The batch size is 50 across all three datasets, and the learning rate schedule follows a cosine decay profile. There are 10 warmup epochs, and each linear probing configuration is trained for 50 epochs.
Table 6: Pre-training settings.
Configuration Value
Optimizer AdamW
Absolute Learning Rate 5e-4
Weight Decay 0.05
Optimizer Momentum $\beta_{1}=0.9,\beta_{2}=0.95$
Batch Size 50
Learning Rate Schedule Cosine Decay
Warmup Epochs 50
Training Epochs
20 (w/ augmentation)
4000 (w/o augmentation)
Table 7: Finetuning settings.
Configuration UCI-HAR WISDM IM-WSHA
Optimizer AdamW AdamW AdamW
Learning Rate 1e-3 (Absolute) 1e-2 (Base) 1e-3 (Base)
Weight Decay 0.05 0.05 0.05
Optimizer Momentum
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
Batch Size 50 64 64
Learning Rate Schedule Cosine Decay Cosine Decay Cosine Decay
Warmup Epochs 5 5 5
Training Epochs 50 50 50
Table 8: Linear probing settings.
Configuration UCI-HAR WISDM IM-WSHA
Optimizer AdamW AdamW AdamW
Learning Rate 1e-3 (Absolute) 1e-1 (Base) 1e-2 (Base)
Weight Decay 0 0 0
Optimizer Momentum
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
$\beta_{1}=0.9,$
$\beta_{2}=0.999$
Batch Size 50 64 64
Learning Rate Schedule Cosine Decay Cosine Decay Cosine Decay
Warmup Epochs 10 10 10
Training Epochs 50 50 50
Algorithm 1 Augment Time Series
  Input: Dataset $\mathcal{X}$, target length $L$.
  Output: Augmented time series $\tilde{x}$.
  Sample $x_{1},x_{2}\sim\mathcal{X}$
  Sample chunk length $\lambda\sim\mathcal{U}(1,L)$ and start indices $s_{1},s_{2}\sim\mathcal{U}(0,L-\lambda)$
  Replace $x_{1}[s_{1}:s_{1}+\lambda]$ with $x_{2}[s_{2}:s_{2}+\lambda]$
  return $\tilde{x}=x_{1}[1:L]$
Appendix E Quantitative analysis
We analyzed the classification results of MoCA using a confusion matrix, which provides detailed insights into the model’s performance by displaying counts of correct and incorrect predictions across different classes. As shown in Figure 6, MoCA demonstrates robust learned representations, effectively capturing patterns in most cases and validating the strength of the pre-training process.
Figure 6: Confusion matrix of classification task using finetuned weights from MoCA Augmentation model, and synchronized masking model on the UCI-HAR test dataset. The rows denote true labels per class, and the columns denote predicted labels. Synchronized masking (left) showed top-1 accuracy of 91.7%, MoCA (middle) showed 96.8% , and MoCA w/o augmentation (right) showed 94.8% for 7-class classification.
(a) Synchronized masking
(b) MoCA w/ augmentation (Ours)
(c) MoCA w/o augmentation (Ours)
Appendix F Visulization
We present visualizations to provide insight into the attention mechanisms and learned feature representations of our model.
Figure 7: t-SNE visualization of UCI-HAR classes as represented by MoCA. For each class, the embeddings are derived from the [CLS] token of all samples in the validation set. MoCA (left) demonstrates distinct separation among classes, particularly between sitting and standing. Synchronize-Masking (mid) shows similar behavior but exhibits entanglement among the walking-up, walking-down, and walking classes. Channel-Mixing (right) presents greater entanglement, with walking and laying classes being less distinct, and standing and laying appear misplaced relative to expected proximities.
Figure 8: Training loss for MoCA w/ aug. over 200,000 steps. The x-axis shows both the number of steps and the corresponding time in hours. The loss rapidly decreases at the start and then continues to decrease gradually, indicating that the model converges as training progresses.
Figure 9: We compute attention scores of the [CLS] token to the patch tokens in the penultimate layers and aggregate them by channel. MoCA demonstrates evenly distributed attention scores across channels, suggesting that each channel contributes to the model’s predictions.